a bus bound for baguio leaves the terminal and averages 75 kph. One hour later,a second bus leaves the same terminal and averages 85 kph. in how many hours will the second bus overtake the first ?
Answer by stanbon(75887)

a bus bound for baguio leaves the terminal and averages 75 kph. One hour later,a second bus leaves the same terminal and averages 85 kph. in how many hours will the second bus overtake the first ?
----
1st bus DATA:
rate = 75 kph ; time = x hrs; distance = 75x km
-------------------------
2nd bus DATA:
rate = 85 kph ; time = x-1 hrs; distance = 85(x-1) km
-------------------------
Equation:
distance = distance
75x = 85x-85
-10x = -85
x = 8.5 hrs (time needed for the 2nd bus to catch up to the 1st bus)
=================
cheers,
Stan H.